OpenStudy (anonymous):

Is this the correct formula: ( = pi*r^2) or this one A=1/2*pi*r^2 @jim_thompson5910 @ganeshie8

OpenStudy (vishweshshrimali5):

1/2*pi*r^2

OpenStudy (anonymous):

With the first equation I got 40 , with the 2nd equation I got 20.

jimthompson5910 (jim_thompson5910):

How did you get 40?

OpenStudy (anonymous):

Check the question title, that's why I'm asking which equation is correct...

jimthompson5910 (jim_thompson5910):

What's the diameter of this semi-circle?

OpenStudy (anonymous):

I already did all that : The two radius' are 5 and 3.5 . I just don't know which formula is correct! If the 1st formula in the title is correct than the answer would be 40, if the 2nd formula in the title is correct than the answer would be 20 instead.

jimthompson5910 (jim_thompson5910):

A = pi*r^2 is the area of a whole circle A = 1/2*pi*r^2 is the area of half a circle

OpenStudy (anonymous):

Ah! That's why.. ------------- I solved it these two ways: 1st equation \[A =\pi*5^2\] \[A= 78.539 \approx 78.54 \] \[A =\pi*3.5^2\] \[A= 38.48 \] \[78.53-38.48=40.05\]

jimthompson5910 (jim_thompson5910):

ok now you can cut that answer in half since you only want the upper half (not the entire circle)

OpenStudy (anonymous):

Now, if you use the second equation it would imply the answer is 20 .

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

:D Yay.